# ADMM Based Semi-Structured Pattern Pruning Framework For Transformer

### v1

@@ Abstract.
Abstract:NLP(natural language processsing) has achieved great success through the transformer model.However, the model has hundreds of millions or billions parameters,which is huge burden for its deployment on personal computer or small scale of server.To deal with it, we either make the model’s weight matrix relatively sparser, or compress attention layer. Pattern pruning ,one of the most important pruning methods, permits selecting fixed number of parameters in each divided pattern block and prunes it. However, the effect of pattern pruning is strictly limited by the sparsity within a region of weights in each layer. In this paper,we first introduced Alternating Direction Method of Multipliers(ADMM) based pattern pruning framework to reshape the distribution of activation map. Specifically, we propose to formulate the pattern pruning on transformer as a constrained optimization and use ADMM to optimize the problem. In this way, the initial dense feature maps is transformed to rather regionally sparsified ones.Therefore, we can then achieve higher compression ratio with better performance based on pattern pruning method. Additionally, this paper provides a theoretical derivations of the ADMM with local sparsity. Finally, we also extend the proposed ADMM based framework on quantization to demonstrate its generalization and use SR-STE to avoid gradient vanishing problem. We conduct extensive experiments on classification tasks over GLUE datasets. Significantly, we achieve 50% percent compression ratio while maintaining 55.4% Matthews correlation on COLA, 68.8% accuracy on RTE and overall score 80.1. Our framework also perform well on other tasks on GLUE datasets.
@@ 4. Evaluation
Implementation. We conduct experiments on NVIDIA 3090 GPU with batch size 24, and use Adamw optimizer(ADAMW, ) with the initial learning rate of 2e-5 meanwhile using learning rate scheduler to make adaptive adjustments during the training process. The code is implemented based on hugging face using python 3.9.3, pytorch 1.13.1 and cuda 12.3.
Former works have indicated that higher sparsity leads to greater performance loss when the pattern size is larger than 4*4. Therefore, for evaluation, we focus on evaluating 50% sparsity, 4*4 pattern size, 32 pattern number since it is applicable in most situations. We use pretrained model checkpoints for BERT2 , provided by the HuggingFace model repository. All models were fine-tuned. Finetuning these models required between 7 minutes (RTE) and 7 hours (MNLI) depending on the scale of the task. We further choose 4*4 pattern size with 80% sparsity and 8*8 pattern size with 50 % sparsity to demonstrate the great effect of the model weight distribution change.
Datasets. We evaluate ADMM-based pattern pruning framework and our baselines using the General Language Understanding Evaluation (GLUE) benchmark(glue, ), a collection of NLP tasks varying in data availability and complexity. We report the Spearman correlation for STS-B, the F1 score for MRPC, Matthews correlation for CoLA, and accuracy for all remaining tasks. The reported average is the geometric mean of reported scores. With GLUE as an evaluation criteria, we can better compare our framework with other pruning methods.
@@ 4.1. Main results
ADMM unstructured:To measure the accuracy cost of semi-structured accuracy specifically, we create another baseline that uses ADMM but induces unstructured sparsity at 50% per-layer (rather than global) sparsity.

### v2

@@ Abstract.
Abstract:NLP(natural language processsing) has achieved great success through the transformer model.However, the model has hundreds of millions or billions parameters,which is huge burden for its deployment on personal computer or small scale of server.To deal with it, we either make the model’s weight matrix relatively sparser, or compress attention layer. Pattern pruning ,one of the most important pruning methods, permits selecting fixed number of parameters in each divided pattern block and prunes it. However, the effect of pattern pruning is strictly limited by the sparsity within a region of weights in each layer. In this paper,we first introduced Alternating Direction Method of Multipliers(ADMM) based pattern pruning framework to reshape the distribution of activation map. Specifically, we propose to formulate the pattern pruning on transformer as a constrained optimization and use ADMM to optimize the problem. In this way, the initial dense feature maps is transformed to rather regionally sparsified ones.Therefore, we can then achieve higher compression ratio with better performance based on pattern pruning method. Additionally, this paper provides a theoretical derivations of the ADMM with local sparsity. Finally, we also extend the proposed ADMM based framework on quantization to demonstrate its generalization and use SR-STE to avoid gradient vanishing problem. We conduct extensive experiments on classification tasks over GLUE datasets. Significantly, we achieve 50% percent compression ratio while maintaining overall score 80.1 on GLUE dataset.
@@ 4. Evaluation
Implementation. We conduct experiments on a NVIDIA 3090 GPU with batch size 24, and use Adamw optimizer(ADAMW, ) with the initial learning rate of 7e-5 meanwhile using learning rate scheduler to make adaptive adjustments during the training process. Both pre-training and pattern pruning were trained for 5 epochs. The experimental section uses Bert as the base model for pruning. In order to test the performance on different downstream tasks, we use Bert-base as an example. The data processing and other experimental settings remain consistent with the default settings of Bert. The experimental environment consists of Python 3.9.3, PyTorch 1.13.1, and CUDA 12.3.
Former work PP-Transformer(PP-Transformer, ) has indicated that higher sparsity leads to greater performance loss when the pattern size is larger than 4*4. Therefore, for evaluation, we focus on evaluating 50% sparsity, 4*4 pattern size, 32 pattern number, as these settings are applicable in most situations. We use pretrained model checkpoints for BERT2 , provided by the HuggingFace model repository. All models require fine-tuning, and the time required for fine-tuning depends on the size of the dataset and tasks, with a maximum duration of 7 hours. We further choose 4*4 pattern size with 50 % sparsity to demonstrate the change in the weight distribution of the model.
Datasets. We evaluate ADMM-based pattern pruning framework and our baselines using the General Language Understanding Evaluation (GLUE) benchmark(glue, ), a collection of NLP tasks varying in data availability and complexity. We report the Spearman correlation for Semantic Textual Similarity Benchmark(STS-B), the F1 score for Microsoft Research Paraphrase Corpus(MRPC), Matthews correlation for the Corpus of Linguistic Acceptability(CoLA), and accuracy for all remaining tasks:MultiNLI(MNLI),The Stanford Sentiment Treebank(SST-2),Question NLI(QNLI),Recognizing Textual Entailment(RTE). The reported average is the geometric mean of reported scores. With GLUE as an evaluation criteria, we can better compare our framework with other pruning methods.
@@ 4.1. Main results
We use Bert as the baseline model. To demonstrate the effectiveness of Pattern pruning (optimized with ADMM), we include ADMM unstructured, N×M Transformer, and Pattern pruning (without ADMM) as comparisons. Here is a detailed explanation of each method:
@@
ADMM unstructured: To measure the accuracy of semi-structured method specifically, we utilize ADMM but introduce unstructured sparsity at a per-layer level instead of applying global sparsity. The objective is to achieve a 50% sparsity rate within each layer.